For the given functions f and g , find the indicated composition.

f(x) = 12x2 - 9x, g(x) = 8x - 10
(f∘g)(6)

Answer by ewatrrr(24785)


Hi
f(x) = 12x^2 - 9x, g(x) = 8x - 10  | g(6) = 38
(f∘g)(6) = 12(38)^2 - 9*38 = 16,986
Wish You the Best in your Studies.